OpenStudy (anonymous):

find the polynomial that has the given zeros.-2, 1, 1, 5

OpenStudy (anonymous):

A polynomial having those as "zeroes" means that when the polynomial function f(x) is evaluated at those points, the function goes to zero. The easiest way to find a function that satisfies this is \[\begin{array}{l} x=-2 & \longrightarrow & x+2=0 \\x=1 & \longrightarrow & x-1=0 \\ x=1 & \longrightarrow & x-1=0 \\x=5 & \longrightarrow & x-5=0 \end{array}\] so \[f(x) = (x+2)(x-1)(x-1)(x-5)\] You can see that plugging in any of the given "zeroes" of the function makes the function 0. Simplify that into a 4th degree polynomial to find your solution ^_^